OpenStudy (anonymous):

Andy's average driving speed for a 4 hour tripwas 45mph. during the first 3 hours he drove 40mph. what was his average speed for the last hour of his trip?

OpenStudy (anonymous):

= 60 mph

OpenStudy (anonymous):

Average speed=total distance/total time total distance=45*4=180,distance in 3 hrs=40*3=120, rem distance=60 which he travelled in 1 hr, average speed=60

OpenStudy (anonymous):

thanks